40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 7 14:51:20 2019, aaron, Update, Electronics, WFS head RF measurements WFShead_response.pdf
    Reply  Mon Oct 7 19:50:33 2019, gautam, Update, IOO, IMC locking not working after this work IMCflaky.pngIMG_8015.JPG
       Reply  Tue Oct 8 10:29:19 2019, gautam, Update, IOO, MC Transmission scan MC2_transmission_scatter.pdftransmissionMaps.pdfcorrelStructure.pdf
       Reply  Tue Oct 8 16:54:56 2019, rana, Update, IOO, IMC locking not working after this work 
    Reply  Tue Oct 8 16:00:06 2019, aaron, Update, Electronics, WFS head RF measurements WFS_ACresponse.pdfWFS_DCresponse.pdf
       Reply  Wed Oct 9 12:15:05 2019, rana, Update, Electronics, WFS head RF measurements 
Message ID: 14952     Entry time: Tue Oct 8 16:54:56 2019     In reply to: 14946
Author: rana 
Type: Update 
Category: IOO 
Subject: IMC locking not working after this work 

I think this offset setting thing is not so good. People do this every few years, but putting offsets in servos means that you cannot maintain a stable alignment when there are changes in the laser power, PMC trans, etc. The better thing is to do the centering of the WFS spots with the unlcoked beam after the control offsets have been offloaded to the suspensions.

ELOG V3.1.3-